SPECTROSCOPIC IMPLICATIONS OF THE “STRUCTURE” OF HELIUM DIHALOGEN COMPLEXES

\underline{SARA E. RAY} and ANNE B. McCOY, Department of Chemistry, The Ohio State University, Columbus, OH 43210; JOHN M. HERBERT, Department of Chemistry, University of California, Berkeley, CA 94720.

The rovibronic (B̃ ← X̃) transitions of He·Br_{2} and He·I_{2} have been studied through calculations of the energies of the bound states on the X and B state potential surfaces. In these studies, literature potentials are used for all but the X-state of He·Br_{2}.^{a}^{b} In that case, we calculated the potential at the CCSD(T) level of theory using the SDD+G(3df) basis for Br, the aug- cc-pVQZ basis for He, and a 3s3p2d2f1g set of bond functions, optimized for rare-gas molecule interactions,^{c} centered at the midpoint between He and the Br_{2} center of mass. Using these potentials, the energies and wave functions of the bound states of the complexes are generated for J ≤ 9. From these results, the vibronic spectra are computed over a range of temperatures from 0.1 to 2.5 K. Excellent agreement with the reported spectra is obtained.

aM. P. de Lara-Castells, A. A. Buchachenko, G. Delgado-Barrio, and P. Villarreal, J. Chem. Phys. \underline{120}, 2182 (2004).

bR.Prosmiti, A. Valdes, P. Villarreal, and G. Delgado-Barrio, J. Chem. Phys. \underline{108}, 6065 (2004).

cS. M. Cybulski and R. R. Toczylowski, J. Chem. Phys. \underline{111}, 10520 (1999).